On 1/1/2014, XYZ signed a 2 year, $6,000 non-interest note and received equipment from ABC Manufacturing. The equipment is estimated to be worth $5,144. Based on XYZ’s credit history, a reasonable market rate of interest on similar loans is estimated to be 9%. Since XYZ has recently obtained similar loans from a local bank, XYZ believes that the market interest rate is considered to be a more reliable estimate. Assume that the equipment will be depreciated over 2 years with no salvage value using the straight-line method. ABC Manufacturing primary business is equipment sales. Therefore, ABC believes that the equipment’s fair value is a more reliable estimate. The cost to manufacture the equipment was $4,400 and ABC carried this equipment as inventory on its books prior to the sale to XYZ. Part 1 - Prepare all required entries for 2014 and 2015. You should create entries for both ABC and XYZ. Assume that both companies amortize discounts using the effective interest rate method. Part 2 - Now assume that both companies amortize discounts using the straight-line interest rate method. Prepare the journal entry for 2014 for ABC and XYZ to record interest.
Answer #1
Part - A
For ABC
Estimate for Amount Borrowed = Fair Value of Equipment = $ 5,144
Date Particulars Debit ($) Credit ($)
1/1/2014 Notes Receivable 6000
        Sales 5144
        Discount on Notes Receivable 856
31/12/2014 Discount on Notes Receivable 412
        Interest 412
(Expn. 1)
31/12/2015 Discount on Notes Receivable 444
        Interest 444
(Expn. 1)
31/12/2015 Cash 6000
        Notes Receivable 6000
Expn. 1
Effective Interest Rate = 8% (Look up factor of 5,144 / 6,000 = .8573)
Interest(Discount Amortization) for 2014 = 5,144 * 8% = 412
Interest(Discount Amortization) for 2015 = (5,144 + 412) * 8% = 444
For XYZ
Estimate for Amount Borrowed = Fair Value based on Market Interest Rate = $ 6,000 * PVF (2,9%) = $ 5,050
Date Particulars Debit ($) Credit ($)
1/1/2014 Equipment 5050
Discount on Notes Payable 950
        Notes Payable 6000
31/12/2014 Interest 455
        Discount on Notes Payable 455
(Expn. 1)
31/12/2014 Depreciation (($ 5,050-0)/2) 2525
        Equipment 2525
31/12/2015 Interest 495
        Discount on Notes Payable 495
(Expn. 1)
31/12/2015 Depreciation (($ 5,050-0)/2) 2525
        Equipment 2525
31/12/2015 Notes Payable 6000
        Cash 6000
Expn. 1
Effective Interest Rate = 9% (as provided that fair value of equipment and notes)
Interest(Discount Amortization) for 2014 = 5,050 * 9% = 455
Interest(Discount Amortization) for 2015 = (5,050 + 455) * 9% = 495
Part - B
For ABC
Date Particulars Debit ($) Credit ($)
31/12/2014 Discount on Notes Receivable 428
        Interest 428
For XYZ
31/12/2014 Interest 475
        Discount on Notes Payable 475
Expn. 2
ABC
Interest = (Total Value - Fair Value)/2 years
Interest = (6000 - 5144)/2
Interest = $ 428
XYZ
Interest = (Total Value - Fair Value)/2 years
Interest = (6000 - 5050)/2
Interest = $ 475